If there are 9 Volts across a 9 Ohm resistor, what is the power in watt

  1. If there are 9 Volts across a 9 Ohm resistor, what is the power in watt
Answer #1

As we know that Power across a resistance of R is = V​​​​​​2 /R = 81/9 = 9 Jule per second = 9 watt